OpenStudy (anonymous):

elect the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (−2, −1)

OpenStudy (madhu.mukherjee.946):

when parallel slopes of 2 lines are equal and when perpendicular slope is - inverse

OpenStudy (madhu.mukherjee.946):

so what do you think the slopes in this case will be

OpenStudy (anonymous):

oppisite?

OpenStudy (madhu.mukherjee.946):

y=x+5 so parallel slope will be 1 and perpendicular slope will be -1

OpenStudy (anonymous):

parallel: y = x − 1 perpendicular: y = −x + 1?

OpenStudy (anonymous):

my options

OpenStudy (madhu.mukherjee.946):

okay ,now to find your answer you gotta use the formula- y-y1=m(x-x1)

OpenStudy (madhu.mukherjee.946):

u will use this formula when you will be asked to find the equation of the line passing through a point or two points

OpenStudy (madhu.mukherjee.946):

so parallel equation will be- y+1=1(x+2) y+1=x+2 y=x+1

OpenStudy (anonymous):

this confused me

OpenStudy (madhu.mukherjee.946):

now perpendicular- y+1=-1(x+2) y=-x-3

OpenStudy (anonymous):

thank you, i hope i'll understand it during my test xD

OpenStudy (madhu.mukherjee.946):

which one say

OpenStudy (anonymous):

?

OpenStudy (madhu.mukherjee.946):

u mention which portion

OpenStudy (anonymous):

d right

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (anonymous):

tyty

OpenStudy (madhu.mukherjee.946):

did you not understand the formula